OpenStudy (accidentalaichan):

Write the equation of the line through (0, −2) and (4, 18) in point-slope form.

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (abb0t):

First step is to find the slope: \(\sf \large \color{red}{\frac{y_1-y_2}{x_2-x_2} = m}\) your next step is to pick ONE point. any point. and plug it into equation gien by @grray7 y\(_1\) = your y-coordinate and you know what your x\(_1\) is, your x-coordinate of THAT POINT that YOU picked. Point slope form, you need to rearrange it in the form: \(\sf \color{green}{y=mx+b}\)